3. Find x2 dx if it exists VI+x3 2
Answer #1

I am attaching the pictures of solution. If you find my answer helpful please press thumbs up button. Thanks

3 2 Ans da Sitx let t = 1+x² - d (t) = d (1+x²) de da d (1) + dx d Tx (x3) - dt dx = 0+32 because ole Lconstant) = 0 m-1 det3 - t = 1+(2) - t = 3 when & t 1+x? (by) El +(00)? because (oo) if kyo k 07 =) +00 because ktoo 8 - t - oo 4 80 S مولا Sitx 2at 3 t [y&④) 3. 9 글 (+)+) +) t) || 3 23, - 2 (GE), 9 (o - 5) (o - 3)1 972 (00) because so - ao-k = k(oo) if k>o u x da 1+x² oo so it is / divergent